Find a and b.

If -3 < x < 3, then a < 1 - 2x < b.

Answer by ikleyn(52786)

.
Find a and b.
If -3 < x < 3, then a < 1 - 2x < b.
~~~~~~~~~~~~~~~~~~~

Function  {1-2x}  is the monotonically decreasing function of x in the interval -3 < x < 3.


Therefore, we should take for "a" the value of the function 1-2x at the right end
of the interval x = 3, i.e.  

    a = 1 - 2*3 = 1 - 6 = -5.


Respectively,  we should take for "b" the value of the function 1-2x at the left end
of the interval x = -3, i.e.  

    b = 1 - 2*(-3) = 1 + 6 = 7.


ANSWER.  a = -5;  b = 7.


CHECK.  We should check if 

    -5 < 1 - 2x < 7  at x in the interval (-3,3).   

        It is correct.

Solved.


////////////////////


Comment from student: You are saying to equate a and b to 1 - 2x and solve individually. Yes?


My response: Hello, you asked me to solve this problem for you.

I did it and expressed it in a way which I found the most appropriate.

Thus I said what I said in my post.

To me, there is no need to re-tell it in other words.